OpenStudy (anonymous):

get the differential from y = x^2 tanh^3 (√x)

sam (.sam.):

\[\huge \frac{dy}{dx}=2xtanh ^{3}\sqrt{x}+\frac{3x ^{\frac{3}{2}}sech ^{2}(\sqrt{x})}{2}\]

OpenStudy (anonymous):

You want us to find the derivative of y = x^2 tanh^3 (√x)?

OpenStudy (anonymous):

Sam's answer looks right some things to note that if you are going to find the derivative where two of the variables are multiplying you have to use the product rule F(x)- x1 * x2 =( x1' * x2 )+ (x2' * x1)